OpenStudy (anonymous):

The spherical head of a straight pin has a radius of 0.8 mm. What is the approximate volume of the head of the straight pin? Use 3.14 to approximate pi and round your answer to two decimal places.

OpenStudy (acxbox22):

the formula for volume is V=(4/3)(pi)(r^3) plug in 0.8 for r 3.14 for pi and solve for V

OpenStudy (anonymous):

I still don't get it

OpenStudy (anonymous):

im sorry im a little slow today \

OpenStudy (igreen):

\(V_{sphere} = \dfrac{4}{3} \pi r^3\) \(V_{sphere} = \dfrac{4}{3} (3.14) (0.8)^3\) Multiply 4/3 * 3.14 * 0.8 * 0.8 * 0.8

OpenStudy (acxbox22):